OpenStudy (anonymous):

A ball is dropped from a height of 21 m and always rebounds of the height of the previous drop. How far does it travel (up and down) before coming to rest?

OpenStudy (anonymous):

you are going to have to sum a geometric series, but you are missing something in the question

OpenStudy (anonymous):

what am i missing?

OpenStudy (anonymous):

read your post

OpenStudy (anonymous):

oh it rebounds 1/3 of the height of the previous drop

OpenStudy (anonymous):

ok now we can do it first off it drops 21 m, put that number aside, we will add in later

OpenStudy (anonymous):

then it goes up 7 and also down 7 as \(21\times \frac{1}{3}=7\) so \[2\times 21\times \frac{1}{3}\] for the first bounce up and down

OpenStudy (anonymous):

second bounce up and down will be \[2\times 21\times \left(\frac{1}{3}\right)^2\] or \[42\times \frac{1}{3^2}\]

OpenStudy (anonymous):

and so on add up \[42\times\frac{1}{3}+42\times \frac{1}{3^2}+42\times \frac{1}{3^2}+...\] or in sigma notation \[42\sum_{n=1}^{\infty}\frac{1}{3^n}\]